A hiker stands on the edge of a cliff 490 metre above the ground and throws the stone horizontally with an initial speed of 15 metre per second.The speed with which it hits the ground is...?

Grade:11

1 Answers

Vikas TU
7 years ago
in horizontal direcn. net accln. = 0
for horzntl. dircn. therfore,
 
Vx  = u + 0 = 15 m/s  …....................(1)
 
For vertical direcn.,
490 = 0 + 0.5gt^2
calculate t from here,
then from first eqn.
Vy = 0 + gt
and get Vy.
 
Resulatant speed = root(Vx^2 + Vy^2)